OpenStudy (anonymous):

determine whether the problem is an expression or an equation. then simplify each expression or solve 3+5(2x-4)-6x

OpenStudy (vishweshshrimali5):

See an equation always carry an equality sign(=) but since there is no c=such sign in the given question

OpenStudy (vishweshshrimali5):

THUS 3 + 5(2x - 4) - 6x is an expression

OpenStudy (anonymous):

how did you figure that?

OpenStudy (vishweshshrimali5):

Now we can simplify the expression in the following way:

OpenStudy (anonymous):

how can I show more detalied work o my paper?

OpenStudy (vishweshshrimali5):

See: 3 + 5(2x - 4) - 6x = 3 + 10x - 20 - 6x = 4x - 17

OpenStudy (vishweshshrimali5):

Therefore , 3 + 5(2x - 4) = 4x - 17

OpenStudy (anonymous):

so theres no answer?

OpenStudy (vishweshshrimali5):

If you are satisfied with my answer then please give me a medal and become my fan

OpenStudy (anonymous):

if its a expression theres no answer right?

OpenStudy (vishweshshrimali5):

Yaa we can never solve an expression we can only simplify it.

OpenStudy (vishweshshrimali5):

Is it clear to you

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

thanx

OpenStudy (vishweshshrimali5):

thanx for giving me a medal